# Booster: Tackling Harmful Fine-tuning for Large Language Models via Attenuating Harmful Perturbation

### v3

@@ Appendix B Baseline Descriptions
  • Vaccine (alignment stage solution). For Vaccine (Huang et al., 2024c), we use Vaccine algorithm at the alignment stage to align the model with the alignment dataset. Then, we use normal SFT to train the model on the user dataset. In our experiment, the hyper-parameter of Vaccine is $\rho=5$ which we select by grid searching over $[0.1,1,2,5,10]$.
  • RepNoise (alignment stage solution). For RepNoise(Rosati et al., 2024a), we use RepNoise algorithm at the alignment stage to align the model with the alignment dataset/harmful dataset. Then, we use normal SFT to train the model on the user dataset. Its hyper-parameter is selected as $\alpha=1$ and $\beta=0.001$.
  • Lisa (fine-tuning stage solution). For Lisa(Huang et al., 2024b), we use SFT at the alignment stage to align the model with the alignment dataset. Then, we use Lisa algorithm to train the model on the user dataset. The regularizer’s intensity of Lisa is set to $\rho=0.01$, which we grid search over $[0.001,0.01,0.1,1]$.
For Booster, as it is an alignment stage solution, we use Booster algorithm (Algorithm 1) to align the model with alignment/harmful dataset. Then we use SFT to finetune the model on user dataset.
Then we introduce the high level idea of each defense baseline.
@@
  • RepNoise (alignment stage solution). Same as Vaccine, RepNoise aims to strengthen the aligned model’s robustness in the alignment stage. The core idea is to degrade the harmful data’s embedding to a random Gaussian noise such that the encoded information is sufficiently destroyed, making it harder to recovered in the later fine-tuning.
@@
  • Lisa (fine-tuning stage solution). Lisa aims to preserve the model alignment knowledge in the fine-tuning stage. The idea is to rotate the optimization into two states. For the first state, the model is optimized over the alignment dataset and for the second state, the model is optimized over the fine-tuning dataset. Further, a proximal term is added to each state’s optimization to guarantee better performance.
Appendix C Booster vs. TAR
A concurrent alignment stage solution TAR (Tamirisa et al., 2024), first released on August 1st (approximately one month before Booster first release) share a similar optimization viewpoint to our proposed Booster solution. We then present a simplified idea of TAR (only one inner step in TAR is taken for sake of easy comparisons with Booster). Notations are also adapted for easy comparison with Booster. Formally, a simplified version of TAR aims to solve the following problem:
where $\tilde{f}(\bm{w})$ is a representation engineering inspired retain loss for preserving performance on the capabilities proxy dataset, and $h(\bm{w})$ is the entropy loss over harmful data. The second term is a tamper-resistance loss that ensures the negative entropy loss over the harmful data to be minimized after one step of harmful fine-tuning attack. Of note, in original TAR, multiple inner steps are taken to simulate the harmful fine-tuning attack.
In constrast, Booster aims to solve the following problem:
The main difference lies in two main aspects.
  • $f(\bm{w})$ vs. $\tilde{f}(\bm{w})$. For optimizing the alignment loss $f(\bm{w})$, Booster need to assume the existence of alignment data. This can ensures the model learns to answer harmful prompt in a refusal way. TAR instead optimizes over the retain loss, which requires the existence of a normal proxy dataset.
  • The main difference lies in the second term. TAR aims to directly minimize the negative entropy loss over harmful data after taking the simulated harmful perturbation. In contrast, Booster aims to minimize the harmful loss reduction rate by minimizing $h(\bm{w})-h(\bm{w}-\alpha\frac{\nabla h(\bm{w})}{\|\nabla h(\bm{w})\|})$. While the difference seems to be subtle, the goal of the two designs, as we mentioned, are different. In our initial attempt, we also experiment on a term similar to tamper resistance loss used by TAR. We minimize over the negative cross entropy loss but the training would become very unstable with this loss, because the model will easily collapse ($h(\bm{w})$ will explode to a large value, and eventually leading the model to constantly repeat a single word for any prompts). Therefore, to guarantee the model’s stability, we eventually adopt the current Booster’s formulation by minimizing the harmful loss reduction rate.
Appendix D Vaccine+Booster
@@
Although Vaccine and Booster are both alignment stage solution, there is still possibility to combine them together to yield better defense performance, considering that their designs are orthogonal to each other. For Vaccine+Booster, we consider to solve such an optimization problem:
@@
where $\hat{f}_{\epsilon}(\bm{w})$ denotes the loss over alignment dataset after applying perturbation $\epsilon$ over the embedding. The first term is the contributed from Vaccine and the second term is from Booster.
To solve such a problem, we follow the paradigm adopted by Vaccine, which basically is to alternatively optimizing over the inner/outer problem. Particularly, for the inner problem, we aim to find the optimal perturbation. Because $\epsilon$ is not presented in the second term, the optimal perturbation is same as Vaccine, as follows:
@@
The problem can be solved by iterative gradient method (e.g., SGD, ADAM,etc) by obtaining its first-order gradient. We next show the detailed Vaccine+Booster algorithm.
@@
8:     $\tilde{g}(\bm{w}_{t})=\tilde{\nabla}f(\bm{w}_{t})\!+\!\lambda\left(\tilde{\nabla}h(\bm{w}_{t})\!-\!\tilde{\nabla}h(\bm{w}_{t}-\alpha\frac{\tilde{\nabla}h(\bm{w}_{t})}{\|\tilde{\nabla}h(\bm{w}_{t})\|})\right)$

### v4

@@ Appendix C Baseline Descriptions
  • Vaccine (alignment stage solution). For Vaccine (Huang et al., 2024e), we use Vaccine algorithm at the alignment stage to align the model with the alignment dataset. Then, we use normal SFT to train the model on the user dataset. In our experiment, the hyper-parameter of Vaccine is $\rho=5$ which we select by grid searching over $[0.1,1,2,5,10]$.
  • RepNoise (alignment stage solution). For RepNoise(Rosati et al., 2024b), we use RepNoise algorithm at the alignment stage to align the model with the alignment dataset/harmful dataset. Then, we use normal SFT to train the model on the user dataset. Its hyper-parameter is selected as $\alpha=1$ and $\beta=0.001$.
  • Lisa (fine-tuning stage solution). For Lisa(Huang et al., 2024d), we use SFT at the alignment stage to align the model with the alignment dataset. Then, we use Lisa algorithm to train the model on the user dataset. The regularizer’s intensity of Lisa is set to $\rho=0.01$, which we grid search over $[0.001,0.01,0.1,1]$.
For Booster, as it is an alignment stage solution, we use Booster algorithm (Algorithm 1) to align the model with the alignment/harmful dataset. Then we use SFT to finetune the model on the user dataset.
Then we introduce the high-level idea of each defense baseline.
@@
  • RepNoise (alignment stage solution). Same as Vaccine, RepNoise aims to strengthen the aligned model’s robustness in the alignment stage. The core idea is to degrade the harmful data’s embedding to a random Gaussian noise such that the encoded information is sufficiently destroyed, making it harder to recover in later fine-tuning.
@@
  • Lisa (fine-tuning stage solution). Lisa aims to preserve the model alignment knowledge in the fine-tuning stage. The idea is to rotate the optimization into two states. For the first state, the model is optimized over the alignment dataset, and for the second state, the model is optimized over the fine-tuning dataset. Further, a proximal term is added to each state’s optimization to guarantee better performance.
Appendix D More experimental results
In our main experiments, we use SST2 for fine-tuning dataset. We next show how different methods perform when fine-tuning on more advanced dataset, e.g., GSM8k. In the following experiments, we use Llama2-7B as base model and GSM8K as fine-tuning task. Other experimental settings are set the same with those in the main experiments.
Harmful ratio $p$. As shown in Table 10, results show that Booster maintain the second smallest average harmful score and the highest fine-tune accuracy among the baselines. Compared to Lisa, the baseline which get the smallest harmful score, Booster obtains a slightly larger harmful score by 1.1, but it gets a significantly higher average fine-tune accuracy by 6.05.
Table 10: Performance analysis for different harmful ratio on GSM8k.
Harmful Score Finetune Accuracy
Methods p=0.05 p=0.1 p=0.15 p=0.2 Average p=0.05 p=0.1 p=0.15 p=0.2 Average
SFT 14.80 23.20 32.40 40.00 27.60 15.20 16.10 16.50 14.40 15.55
Lisa 4.90 10.00 11.70 14.90 10.38 12.80 11.80 12.00 11.40 12.00
Repnoise 16.60 23.80 31.10 34.90 26.60 16.10 13.50 14.60 14.90 14.78
Vaccine 4.50 10.50 18.40 24.60 14.50 13.90 13.60 13.20 13.00 13.43
Booster 7.70 9.50 11.40 17.30 11.48 18.90 18.10 17.50 17.70 18.05
Sample number $n$. As shown in Table 11, the results show that Booster is the second place for harmful score and the first place in maintaining fine-tune accuracy. The baseline that obtains the smallest harmful score is Lisa. Lisa achieves a slightly smaller harmful score by 1.9, while significantly downgrade the finetune accuracy by 5.3 compared to Booster.
Table 11: Performance analysis for different fine-tuning sample number on GSM8k.
Harmful Score Finetune Accuracy
Methods n=500 n=1000 n=1500 n=2000 Average n=500 n=1000 n=1500 n=2000 Average
SFT 10.60 23.20 41.60 58.60 33.50 12.40 16.10 19.00 19.30 16.70
Lisa 4.40 10.00 9.10 9.80 8.33 8.60 11.80 14.40 16.10 12.73
Repnoise 10.00 23.80 36.80 49.30 29.98 11.00 13.50 16.80 19.10 15.10
Vaccine 1.20 10.50 23.60 36.90 18.05 9.90 13.60 16.60 19.10 14.80
Booster 3.60 9.50 18.10 38.50 17.43 13.70 18.10 19.00 21.30 18.03
Base model. As shown in Table 12, the results show that Booster achieves the smallest average HS (harmful score) over three different models, and simultaneously achieve the highest average finetune accuracy (FA).
Table 12: Performance analysis for different models on GSM8k.
Llama2-7B Gemma2-9b Qwen2-7b Average
Methods HS FA HS FA HS FA HS FA
SFT 23.20 16.10 26.40 59.50 37.90 66.80 29.17 47.47
Lisa 10.00 11.80 6.20 54.50 4.40 61.60 6.87 42.63
Repnoise 23.80 13.50 26.20 57.10 25.40 63.70 25.13 44.77
Vaccine 10.50 13.60 18.00 52.50 10.20 63.60 12.90 43.23
Booster 9.50 18.10 2.30 58.40 4.90 70.00 5.57 48.83
Summary. The above results show that Booster is the best at maintaining high fine-tuning downstream task accuracy, while perserving low harmful score compared to the state of the art approaches. We observe that Booster has slightly higher harmful score compared to Lisa (one of the baselins) in some settings, though the Lisa method has a relatively low fine-tune accuracy in most settings. We will show more results to address this concern next.
Appendix E Booster+Lisa
The previous results show that the baseline Lisa achieves a slightly smaller harmful score compared to Booster (though it comes with a large loss of fine-tune accuracy). To merge the benefit of both methods, it is natural to consider to combine the design of Booster and Lisa together. Specifically, Booster is an alignment stage solution, which is operated in the safety alignment stage (before fine-tuning). Lisa is a fine-tuning stage solution, which operates in the fine-tuning stage after safety alignment. They are orthogonal to each other and therefore can be easily combined. Next, we show how the combined design can improve defense performance (all results are produced by GSM8K and Llama2-7B unless otherwise specified).
Harmful ratio $p$. As shown in Figure 13, Booster+Lisa is able to consistently outperform Lisa in terms of harmful score and fine-tune accuracy. Booster+Lisa achieve an astoudning average harmful score of 1.88, while preserving a good average fine-tune accuracy of 13.83.
Table 13: Comparison for Booster+Lisa for different harmful ratio on GSM8k.
Harmful Score Finetune Accuracy
Methods p=0.05 p=0.1 p=0.15 p=0.2 Average p=0.05 p=0.1 p=0.15 p=0.2 Average
Lisa 4.90 10.00 11.70 14.90 10.38 12.80 11.80 12.00 11.40 12.00
Booster 7.70 9.50 11.40 17.30 11.48 18.90 18.10 17.50 17.70 18.05
Booster+Lisa 1.60 2.00 2.10 1.80 1.88 13.60 13.60 14.80 13.30 13.83
Sample number $n$. For different sample number, Booster+Lisa is also able to consistently outperform Lisa in both two metrics in all groups of experiments. The harmful score for Booster+Lisa is extremely small (only a 1.88 average harmful score is reported).
Table 14: Performance analysis for different fine-tuning sample number on GSM8k.
Harmful Score Finetune Accuracy
Methods n=500 n=1000 n=1500 n=2000 Average n=500 n=1000 n=1500 n=2000 Average
Lisa 4.40 10.00 9.10 9.80 8.33 8.60 11.80 14.40 16.10 12.73
Booster 3.60 9.50 18.10 38.50 17.43 13.70 18.10 19.00 21.30 18.03
Booster+Lisa 1.40 2.00 2.10 2.00 1.88 9.70 13.60 15.40 16.10 13.70
Base model. The same advantage is observed for Booster+Lisa by testing on different models. Particularly, we show that the performanace of Booster+Lisa is even more pronounced for more advanced model Gemma2-9B and Qwen2-7B. For example, for Gemma2-9B, Booster+Lisa achieves a smaller harmful socre by 2.7, and a remarkable boost of fine-tune accuracy by 7.5 compared to Lisa.
Table 15: Performance analysis for different models on GSM8k.
Llama2-7B Gemma2-9b Qwen2-7b Average
Methods HS FA HS FA HS FA HS FA
Lisa 10.00 11.80 6.20 54.50 4.40 61.60 6.87 42.63
Booster 9.50 18.10 2.30 58.40 4.90 70.00 5.57 48.83
Booster+Lisa 2.00 13.60 1.10 61.10 1.70 69.10 1.60 47.93
Summary. While Booster obtains a slightly higher harmful score compared to Lisa, the combined design Booster+Lisa outperforms Lisa in all groupds of experiments in terms of maintaining even smaller harmful score and significantly higher fine-tune accuracy. This remarkable result cannot be obtained without our Booster design. Given this, the contribution of our Booster should not be downplayed.
Appendix F Vaccine+Booster
@@
Although Vaccine and Booster are both alignment stage solutions, there is still a possibility to combine them together to yield better defense performance, considering that their designs are orthogonal to each other. For Vaccine+Booster, we consider to solve such an optimization problem:
@@
where $\hat{f}_{\epsilon}(\bm{w})$ denotes the loss over alignment dataset after applying perturbation $\epsilon$ over the embedding. The first term is contributed from Vaccine and the second term is from Booster.
To solve such a problem, we follow the paradigm adopted by Vaccine, which basically is to alternatively optimize over the inner/outer problem. Particularly, for the inner problem, we aim to find the optimal perturbation. Because $\epsilon$ is not presented in the second term, the optimal perturbation is the same as Vaccine, as follows:
@@
The problem can be solved by the iterative gradient method (e.g., SGD, ADAM, etc) by obtaining its first-order gradient. We next show the detailed Vaccine+Booster algorithm.
@@
8:     $\tilde{g}(\bm{w}_{t})=\tilde{\nabla}\tilde{f}_{\epsilon^{*}}(\bm{w}_{t})\!+\!\lambda\left(\tilde{\nabla}h(\bm{w}_{t})\!-\!\tilde{\nabla}h(\bm{w}_{t}-\alpha\frac{\tilde{\nabla}h(\bm{w}_{t})}{\|\tilde{\nabla}h(\bm{w}_{t})\|})\right)$
@@
Appendix G More insights from the observation
G.1 Alternative insights
By Figure 2, we in Section 3.2 systematically study the impact of harmful perturbation and motivate our idea to reduce the harmful loss reduction rate over harmful data. Because in the real fine-tuning stage, the model is fine-tuned on a mixture of fine-tune data and harmful data. A natural question is that can we utilize the fine-tune data for defense design? We repeat the same figure here for illustration of this alternative idea.
Figure 4: Model statistics (Left: harmful score, Middle: harmful training loss, Right: harmful testing loss) after fine-tuning on pure SST2/harmful data for different steps. Specially, harmful training loss refers to the loss over the harmful data used in training, while harmful testing loss refers to that over the testing harmful data that the model never sees in fine-tuning stage.
Alternative Insights. Let’s assume the fine-tuning dataset is a SST2 dataset and in the real fine-tuning stage, the model is fine-tuned on a mixture of SST2 data and harmful data. We already observe that fine-tuning on harmful data will lead to a drastic reduction on harmful loss, and therefore resulting in an increase of harmful score of the model. However, we also observe that fine-tuning on SST2 data slightly increase the harmful loss, though it does not significantly alter the harmful score. An possible idea for optimization is that maybe we can optimize the aligned model such that when it is fine-tuned on SST2 data, the harmful loss can be drastically increased (i.e., the harmful loss increase ratio is high when fine-tuing on benign fine-tune data). By taking steps on the benign data (e.g., SST2), we could possibly balance out the harmful loss reduction when the model is taking steps on harmful data. This intuitive idea could be a possible direction of future defense design. We leave this idea for our future work, but we encourage interested readers to pursuit as well.
G.2 A Strange Phenomenon
By Figure 3 (or Figure 5 repeats in the following), we show that the reduction curve of harmful training loss of Booster is smoother, i.e., the harmful loss reduction rate is smaller, which perfectly reflects Booster’s optimization objective. However, we also observe a strange phenomenon in the middle of Figure 3 that the harmful training loss of booster at step 0 is lower than that of SFT. We in this section provide more information on the reasons leading to this phenomenon.
Figure 5: Model Statistics (Left: harmful score, Middle: harmful training loss, Right: harmful testing loss) after fine-tuning on 10% of harmful data for different steps. Specially, harmful training loss refers to the loss over the harmful data used in fine-tuning, while harmful testing loss refers to that over the testing harmful data which the model never see in fine-tuning phase.
More information. To understand why the initial point (i.e., the aligned model) in the fine-tuning stage has a lower harmful training loss, we need to looks closer on how this initial point (i.e., the aligned model) is formed in the alignment stage. We next show in Table 16 how the alignment loss/harmful training loss evolves when the model is aligned by SFT and Booster. As shown, with the alignment steps increase, the alignment loss of both SFT and Booster are decreased. However, the major difference is how harmful loss is changing with alignment steps. The harmful loss of SFT is increased to a large value (from 1.582 to 3.920), while Booster’s harmful loss is only increased slightly from 1.582 to 1.736. This result explains why in the above figure, Booster has a relatively low initial harmful training loss while SFT has a relatively high loss.
Table 16: Alignment loss/harmful loss with alignment steps. Alignment loss means the loss over the alignment data and harmful loss means that loss over the harmful data used for Booster’s alignment.
Alignment steps 0 1000 2000 4000 6000 8000 10000
SFT (Alignment loss) 1.284 0.308 0.119 0.056 0.024 0.021 0.013
SFT (Harmful loss) 1.582 2.287 3.017 4.227 4.415 4.191 3.920
Booster (Alignment loss) 1.284 0.323 0.146 0.053 0.042 0.023 0.0194
Booster (Harmful loss) 1.582 1.572 1.582 1.638 1.770 1.699 1.736
Conjecture. The result seems to indicate that optimizing the Booster’s regularizer has some influence over the generalization between safety alignment loss and harmful loss. In other words, if we want to make the harmful loss landscape of the model smoother by the Booster’s regularizer, the generalization property between safety alignment loss and harmful loss will also be changed.
G.3 More measurement Metrics
In the main experiment, we test the fine-tuned model performance after fine-tuning on downstream tasks. Reporting only this result raises concern on whether the model aligned by an alignment stage solution (e.g., Booster) have performance degradation on general reasoning capability. To address this concern, we show in Table 17 how the harmful score and the benchmark accuracy differs for different alignment methods.
Table 17: Harmful score/benchmark accuracy measured on aligned model. The benchmark accuracy is the accuracy for GSM8K. Of note, different from fine-tune accuracy, this metric is tested on the aligned model before fine-tuning. The base model we use is Gemma2-9B and other setting are default.
Harmful Score Benchmark Accuracy
SFT 1.5 13.8
RepNoise 0.9 12.8
Vaccine 0.6 9.5
Booster 0.7 13.4
As shown, compared to SFT, Booster slightly decreases the model’s benchmark accuracy by 0.4%. By contrast, the other two alignment stage defenses RepNoise and Vaccine respectively decrease the accuracy by 1% and 4.3%. All the three alignment methods slightly reduce the harmful score of the aligned model. This result indicates that the proposed Booster solution will not significantly hurt aligned LLM’s general reasoning performance.
Appendix H Limitations and Future Extension
The proposed Booster method has a limitation due to its design. Particularly, the exact setting of the hyper-parameters, e.g., the regularizer penalty $\lambda$ and the inner step size $\alpha$, are important for guaranteeing defense and downstream task’s performance. We observe that the exact optimal setting of the hyper-parameters differ from different downstream tasks. However, once the model is aligned by Booster, the model should be able to applied to many unknown downstream tasks. This means that the service provider cannot tune the optimal hyper-parameters for each specific downstream task, but there should be one set of hyper-parameters that work for any downstream task. Finding such set of hyper-parameter is challenging for Booster and also it may not achieve optimal performance for each downstream task, potentially causing trouble in the deployment process. We in this section would like to point out this issue. However, to ensure fair evaluation, in our experiment we do make effort to fix the same set of hyper-parameters for evaluating all downstream tasks (see Table 3).
As evidenced by Ye et al. (2024); Li et al. (2024b), harmful fine-tuning attack might pose a more serious threat in federated instruction fine-tuning, and cannot be mitigated by existing defenses on Byzantine/backdoor attack. To this end, it is interesting to extend the proposed Booster idea to federated instruction fine-tuning context. Techniques such as sparse training (Huang et al., 2024b; 2022; 2023) and quantization (Li et al., 2023b) might be able to be applied here. Another idea is to extend harmful fine-tuning attack/defense to LLM agent research (Wang et al., 2024b; Zheng et al., 2024; Hu et al., 2023; 2024), which represent a more realistic scenario that the attack might pose negative effect.